OpenStudy (anonymous):

F(x)= 1/(x+5) Then f(x+h)-f(x)/h = A/(x+B)(x+Ch+D) A=-1 B=5 C=1 Find D

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ x+5 }\]

OpenStudy (anonymous):

\[\frac{ f(x+h)-f(x) }{ h }= \frac{ A }{ (x+B)(x+Ch+D) }\]

OpenStudy (anonymous):

\[\frac{ -1 }{ x^2+10x+hx+h^2+25 }\]

OpenStudy (anonymous):

A=-1 B=5 C=1 I can't figure out D